OpenStudy (anonymous):

Factor the expression. k2 – 9h2

OpenStudy (anonymous):

bionmial \[(a+b)(a-b) = a^2 - b^2\]

OpenStudy (anonymous):

I don't get it do we plug them in

OpenStudy (anonymous):

formula is actually called difference of squares, and you can also have numbers example: \[(a+4b)(a-4b) = a^2 -4ab +4ab -16b^2\]

OpenStudy (anonymous):

this uses FOIL, and -4ab +4ab cancel out

OpenStudy (anonymous):

can you try the same with the one in the problem? \[k^2 - 9 h^2\]

OpenStudy (anonymous):

\[(a+4b)(a-4b) = a^2 -16b^2\] similar: \[(k+..)(k-..) = k^2 -9h^2\]

OpenStudy (anonymous):

how do I do this again I am confused

OpenStudy (anonymous):

The original equation looks like this \[k^2 - 9 h^2\]

OpenStudy (anonymous):

so it looks like a difference of squares

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the formula for factoring a difference of squares: \[(a+b)(a-b) = a^2 - b^2\]

OpenStudy (anonymous):

what do I plug in for a and b

OpenStudy (anonymous):

k and h